### v1

# The X-ray properties and structure of A3571 up to R500

11email: yuheng@bnu.edu.cn 22institutetext: State Key Laboratory of Particle Astrophysics, Institute of High Energy Physics, Chinese Academy of Sciences, Beijing 100049, China
22email: jiasm@ihep.ac.cn 33institutetext: Max Planck Institute for Extraterrestrial Physics, Giessenbachstrasse 1, 85748 Garching, Germany 44institutetext: Institute for Frontiers in Astronomy and Astrophysics, Beijing Normal University, Beijing 102206, China 55institutetext: National Astronomical Observatories, Chinese Academy of Sciences, 20A Datun Road, Beijing 100101, China 66institutetext: Department of Astronomy, Xiamen University, Xiamen 361005, China 77institutetext: School of Astronomy and Space Science, University of Chinese Academy of Sciences, 19A Yuquan Road, Beijing 100049, China
The X-ray properties and structure of A3571 up to R500
Abell 3571 is a nearby, X-ray bright galaxy cluster located in the Shapley Supercluster. Although it appears morphologically relaxed in X-ray images, multiwavelength observations reveal subtle indications of residual dynamical activity, likely associated with past merger events. Using wide-field () data from the Einstein Probe Follow-up X-ray Telescope (EP-FXT), we extend measurements of the cluster’s properties beyond its radius. We detect surface-brightness excesses on both the northern and southern sides, consistent with previous XMM-Newton results. The temperature, pressure, and entropy in the northern excess region are lower than the average values, whereas those on the southern side are slightly higher. However, we find no evidence for cold fronts or shocks. These features can be interpreted as sloshing motions triggered by an off-center minor merger. Our findings suggest that, despite its symmetric appearance, A3571 is still recovering from a minor merger and is currently in a post-merger phase. This work also demonstrates the excellent capability of EP-FXT for probing the outskirts of galaxy clusters.
Key Words.:
Galaxy clusters — X-ray — Abell 3571 — Intracluster medium — merger dynamics1 Introduction
Galaxy clusters are the most massive gravitationally bound systems in the universe. The majority of their baryonic matter exists in the form of a hot, diffuse intracluster medium (ICM), which emits strongly in the X-ray band (Kravtsov and Borgani 2012; Vikhlinin et al. 2014). Although many clusters appear to be approximately relaxed and in equilibrium in observations, their formation is fundamentally a highly dynamic process. Frequent merger events and continuous matter accretion produce pronounced dynamical effects in the ICM, driving significant departures from hydrostatic and thermal equilibrium (Markevitch and Vikhlinin 2007; Kravtsov and Borgani 2012; Rahaman et al. 2022; Liu et al. 2018). Understanding these disturbances is therefore essential for constraining the physical processes that shape the formation and evolution of galaxy clusters, as well as for providing key observational insights into the assembly of large-scale cosmic structures.
X-ray observations have revealed the characteristic signatures of such dynamical activity, confirming the widespread presence of non-equilibrium structures in galaxy clusters, among which shocks and cold fronts are the most representative. Shocks, typically generated by high-velocity mergers, heat the ICM on short timescales and appear in X-ray images as sharp, arc-shaped edges(Markevitch and Vikhlinin 2007). The Bullet Cluster (Koda et al. 2008; Markevitch et al. 2002) and Abell 3667 (Owers et al. 2009) provide two of the most striking examples. Cold fronts, by contrast, appear as contact discontinuities where denser but cooler gas is juxtaposed against hotter surroundings(Markevitch et al. 2000; Vikhlinin et al. 2001; Markevitch and Vikhlinin 2007). They are generally linked to the displacement of cool core gas induced by gravitational perturbations, such as those triggered by minor mergers (Ascasibar and Markevitch 2006; Roediger et al. 2011; ZuHone et al. 2010). In relaxed clusters, such perturbations can drive sloshing motions of the central gas, producing multiple cold fronts at different radii. Over time, these sloshing motions may evolve into intricate spiral structures, which are observed as concentric edges in X-ray images (Sonkamble et al. 2024; Kadam et al. 2024a, b). To date, spiral-like cold front structures have been identified in a number of well-studied clusters, including Abell 496 (Ghizzardi et al. 2014), Abell 795 (Kadam et al. 2024a), Perseus (Walker et al. 2017; Fabian et al. 2006a), and Virgo (Roediger et al. 2011; Zheng et al. 2025).
Abell 3571 (A3571) is a nearby rich galaxy cluster at redshift 0.039 (Rossetti and Molendi 2010), located in the Shapley Supercluster (Raychaudhury et al. 1991). The brightest cluster galaxy (BCG) of A3571 is MCG–05–33–002, which exhibits a pronounced north–south elongation extending over several hundred kpc (Kemp and Meaburn 1991). Although A3571 is classified as a relaxed cool-core cluster, Laganá et al. (2019) found no evidence of a central temperature drop in its temperature map. In addition, Hudaverdi et al. (2005) reported a noticeable north-south temperature asymmetry, suggesting that A3571 may have experienced recent or ongoing dynamical evolution.
Early X-ray observations revealed that the ICM of A3571 exhibited an average temperature of keV from EXOSAT observations (Edge et al. 1990) and keV based on ASCA data (Markevitch et al. 1998) (the radial temperature profile extends to at ASCA’s half-power diameter (HPD)), with a central metal abundance reaching as high as 0.6 times the solar value at R < and declining to about 0.3 at a radius of , while the XMM-Newton radial profiles further trace the temperature and metallicity distributions out to approximately (Hudaverdi et al. 2005). ROSAT imaging revealed a nearly spherical X-ray brightness distribution, with the intracluster gas appearing azimuthally symmetric, exhibiting only slight ellipticity (Nevalainen et al. 2000). Based on high-resolution observations from XMM-Newton, the central region of A3571 shows pronounced temperature and metallicity inhomogeneities, which appear as a patchy distribution. A temperature difference of up to 2 keV is detected between the northern and southern regions, suggesting that the cluster may have undergone complex dynamical processes, such as the infall of subgroups (Hudaverdi et al. 2005).
Studies of A3571 not only examined the properties of the hot gas but also attempted to infer the gravitational mass and dark matter distribution within the cluster. Nevalainen et al. (2000) combined ASCA temperature data with ROSAT imaging data to estimated the total mass of the cluster to be approximately with Mpc at 90% confidence, which the cluster is in hydrostatic equilibrium. Subsequently, Nevalainen et al. (2001) derived the total mass of A3571, , by combining the BeppoSAX and ASCA temperature profiles with a dark matter density model under the assumption of hydrostatic equilibrium. The corresponding gas mass fraction was estimated to be .
Although X-ray observations indicate that A3571 exhibits an overall spherically symmetric structure with a central cooling flow, multiwavelength data have revealed a more complex dynamical history. Early studies found that the spatial distribution of galaxies and the pronounced north–south elongation of the central cD galaxy are prominent characteristics (Kemp and Meaburn 1991), possibly suggesting that the cluster underwent merger or disturbance events. Subsequently, spectroscopic analyses revealed an asymmetric radial velocity distribution among member galaxies and identified multiple velocity substructures, further supporting an evolutionary scenario in which several smaller galaxy groups may have successively merged into the main cluster (Quintana and De Souza 1993). Venturi et al. (2002) conducted a radio analysis of the galaxy cluster complex consisting of A3571, A3572, and A3575, and concluded that the system is in a late merger stage, with its radio properties reflecting a relatively relaxed dynamical state. More recent multiwavelength observations have not detected large-scale radio halos or significant AGN activity, indicating that A3571 is currently in a relatively relaxed, post-merger evolutionary stage.
Early multiwavelength observations suggest that A3571 is in a relatively relaxed, late post-merger stage. However, our understanding of its detailed merger history remains limited, particularly regarding the hot gas in the outskirts. Therefore, we aim to conduct a more systematic and in-depth investigation of the gas distribution and merger state of A3571 using large field-of-view (FoV) X-ray imaging observations.
EP-FXT is one of the key payloads aboard the Einstein Probe (EP) mission, which was launched on January 9, 2024 (Yuan et al. 2022, 2025). It features a FoV of approximately 1 square degree, an angular resolution of per pixel, and a spatial resolution of in terms of HPD, with an energy band of 0.3-10 keV. The EP-FXT also benefits from an exceptionally low particle background, making it well suited for detecting faint diffuse X-ray emission.
This paper is organised as follows; in Sect. 2, we describe the data used in our analysis. In Sect. 3, the analysis of the X-ray and optical observations is described. In Sect. 4, we discuss cluster dynamics and possible subclusters. We summarise our results in Sect. 5. We adopt (Planck Collaboration et al. 2020), and the best-fit parameters from the spectral fitting are reported with 90% confidence intervals.
2 Observations and data reduction
2.1 X-ray data
During the Performance Verification (PV) phase of the Einstein Probe mission, A3571 was observed with the EP-FXT in full-frame (FF) mode from March 30 to 31, 2024 (ObsID 11904194436), with the total effective exposure time of 40ks. The data reduction followed standard procedures using the Follow-up X-ray Telescope Data Analysis Software (FXTDAS, version 1.20)111http://epfxt.ihep.ac.cn/analysis (Zhao et al. 2025), with additional support from general analysis tools provided by CIAO (version 4.16)222https://cxc.cfa.harvard.edu/ciao/ and HEASoft (version 6.33.1)333https://heasarc.gsfc.nasa.gov/docs/software/lheasoft.
We used the fxtchain pipeline to generate a series of calibrated event files for subsequent scientific analysis. To eliminate the effect of out-of-time (OoT) events, which account for approximately 0.23% of the total detected events in the FF mode, simulated OoT event files were created using fxtootest and were subsequently subtracted from the original images using ftimgcalc. Since the particle background is uniformly distributed and not affected by the telescope’s vignetting function, it was removed before applying vignetting correction. The background spectrum was extracted from the storage area of the event file and modeled using fxtbkggen (Zhang et al. 2025), after which it was subtracted from the image. Vignetting, caused by the decrease in effective area at larger off-axis angles, leads to reduced photon flux at the edges of the FoV. To correct for this effect and improve the SNR, we used reproject_image from CIAO to co-align and merge the images and vignetting-corrected exposure maps from FXT-A and FXT-B. The merged images were then divided by the combined exposure map using farith to produce a vignetting-corrected image. After vignetting correction, point sources were identified using the wavdetect from CIAO and excluded to avoid contamination in the analysis of diffuse emission. Subsequently, the smooth were conducted with a Gaussian kernel. The final X-ray image in the 0.3–7.0 keV band is presented in Fig. 1.
To compare EP-FXT with XMM-Newton, we selected the EPIC-MOS1 observation of A3571 conducted by XMM-Newton in July 2002 (ObsID 0086950201, with an exposure time of 33 ks) and processed the data using the XMM-Newton Science Analysis System (SAS 444https://www.cosmos.esa.int/web/xmm-newton/sas) in FF mode. Time intervals affected by flares were removed by screening the high-energy light curve.
2.2 optical data
The photometric catalog used in this study was derived from the DESI Legacy Imaging Surveys DR10 (Dey et al. 2019). We select a field encompassing the A3571 cluster and its surrounding region, spanning and , with a magnitude limit of magr 22 . Cluster member candidates are identified using the red sequence method, by selecting galaxies within 3 of the best-fit red sequence and further restricting their photometric redshifts to the range . This yields a sample of 399 galaxies.
To examine the spatial distribution of galaxies in and around A3571, we generate a smoothed luminosity density map based on this photometric sample. Following the method of Wen and Han (2013), the r-band luminosity of each galaxy is estimated using its redshift and apparent magnitude. These luminosities are then mapped onto a 200 200 grid covering the selected field, and a two-dimensional (2D) Gaussian kernel with a full width of is applied to smooth the distribution.
The resulting luminosity density map is shown in Fig. 2. A prominent central peak (the yellow triangle) is evident, coinciding with the location of the BCG (marked by a red cross) very well. The galaxy distribution exhibits a mild elongation along the north–south direction. The EP-FXT field (represented by a green square) fully covers both the core and the extended outskirts of the cluster. The overlaid EP-FXT X-ray isophotes (shown as black lines) show a more relaxed morphology. Considering that the relaxation time of the gas is shorter than that of the galaxies, this means that this cluster might be a late-stage merger, where the gases has fully merged and the galaxies have not yet relaxed.
There is another cluster A3572 located about to the south of A3571. Struble and Rood (1999) estimated a redshift of 0.052 and a velocity dispersion of 1216 km s-1 with 6 galaxies. While Lauer et al. (2014) measured a redshift of based on an elliptical galaxy thought to be its BCG. Based on the photometric luminosity density map and the corresponding X-ray emission distribution, no significant enhancement in galaxy number density or X-ray surface brightness is observed at the location of A3572. Thus we conclude that A3572 is not a real cluster.
3 ANALYSIS and RESULTS
The galaxy density distribution of A3571 shows a noticeable elongation along the north–south direction, suggesting that the cluster may not be in a state of hydrostatic equilibrium. To investigate its thermodynamic properties and assess potential merger activity, we conducted a detailed imaging and spectral analysis of the ICM.
3.1 1D analysis
3.1.1 The Surface brightness profile
To comprehensively assess the background level of EP-FXT, we extracted the radial surface brightness profile of A3571, centered on its X-ray emission peak, and compared it with the corresponding profile obtained from XMM-Newton observations, as shown in Fig. 3.
Black points are the EP-FXT data, which extends out to approximately . The blue horizontal line represents the EP-FXT particle background cts/s/arcmin2 , generated using fxtbkggen as described by Zhang et al. (2022). The green horizontal line represents the EP-FXT sky background , obtained by fitting a constant to the surface brightness in the – region of A3571, which nevertheless includes faint emission from the cluster. The blue points indicate the EP-FXT surface brightness after subtracting both background components. We employed a double model to fit the background-subtracted surface brightness profile. The model equation is given below:
The best-fit parameters are listed in Table 1. Additionally, we plotted the two components of the double model, represented by the green and orange dashed lines, which reveal the presence of a dense core in the central region of A3571.
Similarly, we performed a radial analysis of the XMM-Newton MOS1 data and normalized its surface brightness profile to the maximum value of the EP-FXT data. The background components included the sky background obtained from blank-sky observations, and the particle background derived from the 10 – 12 keV spectrum. The cyan points represent the normalized XMM-Newton data, while the blue and green dashed lines correspond to the normalized particle background and the normalized sky background , respectively.
As shown in Fig. 3, the surface brightness profiles of EP-FXT and XMM-Newton are consistent. However, at a radius of around , the XMM-Newton surface brightness profile appears slightly higher. This minor difference is likely attributable to increased background uncertainties and reduced effective exposure toward the edge of the XMM-Newton FoV. In addition, EP-FXT offers a much larger FoV, extending to approximately . Additionally, the low particle background of EP-FXT provides an ideal condition to study diffuse sources. On the other hand, the sky background of EP-FXT is slightly higher than that of XMM-Newton, because it is estimated by fitting a constant to the surface brightness in the outer region of A3571, which still contains faint emission from the cluster. In contrast, the XMM-Newton background is based on blank-sky observations and is therefore free from cluster contamination.
| Parameter | Value |
Note. Best-fit parameters of the double model with errors, based on the background-subtracted. and are in units of arcmin. A, B and C are in units of .
3.1.2 Global spectral analysis
In the spectral analysis, we considered both the particle background and the sky background components. The particle background was generated using the fxtbkggen tool, while the sky background was extracted in a circle region beyond (see the white circle in Fig. 1). The source spectrum was obtained from a circular region with a radius of centered on the X-ray peak (indicated by the red and blue points in Fig. 4). For background, we first subtracted the particle background—scaled to match the sky background region—from the spectrum of the background region shown in Fig. 1.The sky background was then rescaled to match the area of the source region. The sky background and background-subtracted source spectrum are shown as the purple and orange points in Fig. 4, respectively. Similarly, since the particle background generated by fxtbkggen covers the entire CCD, it also needed to be rescaled to match the area of the source region before subtraction. The particle background is shown as the green and brown points in Fig. 4.
The particle background spectra of both FXT-A and FXT-B exhibit similar shapes, remaining at a level of approximately counts s-1 keV-1, which is well below the source emission. Three prominent instrumental lines are present: Al-K (1.48 keV), Fe-K (6.38 keV), and Cu-K (8.04 keV). Around 1 keV, where the source emission peaks, the particle background contributes less than 1% of the source count rate, thereby enabling more reliable analysis of faint diffuse sources. Below 1 keV, the FXT-B spectrum differs from that of FXT-A, primarily due to FXT-B’s larger effective area in this energy band. In addition, extended sources are influenced by atmospheric absorption at low energies because of the satellite’s low-Earth orbit. Therefore, we restrict our data processing and analysis to the 1.0–8.0 keV range.
3.1.3 The radial temperature profile
Thanks to the large FoV of EP-FXT, we are able to explore the outer ICM of A3571 efficiently. Centering on the X-ray peak, we used XSPEC (version 12.14.1) to extract spectra from 11 concentric annular regions (more than 20000 photons were detected in each region) in the 1.0–8.0 keV energy band and conducted a radial analysis. The Ancillary Response Files (ARFs) were generated using the EP-FXT response tool fxtarfgen. The spectra were grouped using grppha with a minimum of 30 counts per bin to ensure the applicability of the statistic. The diffuse ICM thermal emission was modeled using the APEC (version 3.0.9), while Galactic hydrogen absorption was accounted for with the TBABS model. We adopted the solar abundance table from Asplund et al. (2009), with the redshift fixed at 0.039, the abundance fixed at 0.3 and the hydrogen column density fixed at (Willingale et al. 2013). All other parameters were left free during the fit. The radial profiles are shown in Fig. 5. The black points represent the EP-FXT data in this work, respectively. The blue points indicate the XMM-Newton data, while the green points show the ASCA data from Hudaverdi et al. (2005).
In the central region (), the EP-FXT temperature is consistent with the results of Hudaverdi et al. (2005) based on XMM-Newton and ASCA, thereby reinforcing the evidence for the presence of a cool core. Beyond the central region, the temperature profile is broadly consistent with an approximately isothermal distribution at keV. A mild decrease in temperature is observed around , followed by a slight recovery at larger radii.
To further investigate the systematic deviation between temperature of EP-FXT and measurements from literature, we divided the region into northern () and southern () sectors with west as and angles increasing clockwise. The corresponding temperature distributions are shown as the yellow and purple points in Fig. 5.
Although the azimuthally averaged temperatures measured by EP-FXT are overall lower than those obtained from XMM-Newton and ASCA, with a difference on the order of 1 keV, the temperature measured by EP-FXT in the southern region is higher than in the other two datasets. This behavior may reflect the presence of multiple temperature components within the cluster, to which different instruments respond differently, thereby leading to discrepancies in the derived temperatures; nevertheless, the overall temperature trends remain consistent.
Notably, the northern region also exhibits a clear temperature decline. This feature is associated with the brightness excess detected in the northern region and will be explored further in the following sections.
3.2 2D analysis
3.2.1 The brightness residual image
To investigate the spatial structure of A3571, we used Sherpa version 4.16.0555https://cxc.cfa.harvard.edu/sherpa/ to subtract an elliptical -model combined with a constant sky background value (see Sect. 3.1.1) from the Fig. 1. The best-fit parameters of the elliptical model are listed in Table 2, and the residual map is shown in Fig. 6.
A distinct U-shaped surface brightness deficit is found south of the cluster center, whereas a fan-shaped brightness enhancement is detected to the north. Hudaverdi et al. (2005) reported similar features based on XMM-Newton observations. There is also a faint surface brightness enhancement in the southwestern. These features suggest that A3571 is not fully relaxed and still being disturbed.
| Parameter | Value | Unit |
| pixel | ||
| Ellipticity | – | |
| rad | ||
| Amplitude | counts s-1 pix-1 | |
3.2.2 thermodynamic 2D maps
To enable high-resolution 2D mapping of the ICM thermodynamic properties, we performed an Voronoi tessellation on the X-ray image, growing regions to SNR 60 and refining them with SNR‐weighted Lloyd iterations, yielding 102 adaptive bins. The spectrum extracted from each region was modeled with a TBABS * APEC model in 1.0-8.0 keV, with the same parameter settings described in Sect. 3.1.3. Based on the best-fitting temperature and normalization values, we derived pseudo-pressure and pseudo-entropy maps using and , respectively (Sasaki et al. 2016), is electron density and kT is temperature. Because the normalization obtained from spectral fitting is proportional to the square of the electron density, we express and (Fabian et al. 2006b) (see Fig. 7).
The temperature map results indicate that the northern region is overall cooler than the southern region, a trend that is not only consistent with previous XMM-Newton studies (Hudaverdi et al. 2005) but also in agreement with the temperature profile shown in Fig. 5. In the overall cooler northern region, the fan-shaped structure shows an even lower temperature, which is consistent with the drop in temperature around in Fig. 5, supporting the interpretation that this decrease is driven by the fan-shaped structure. In addition, the surface brightness enhancement found in the southwest coincides with a region of higher temperature.
The pseudo-pressure map clearly indicates that the pressure is lower in the northern region and higher in the southern region. In contrast, the pressure distribution derived from XMM-Newton observations appears smooth, showing no prominent localized features, and declines steadily with radius in a similar manner along different azimuthal directions (Hudaverdi et al. 2005). This difference may arise from the different methodologies used to estimate temperature and density. In the case of Hudaverdi et al. (2005), the temperature was estimated from hardness ratios, and the electron density was inferred from the soft-band (0.8–1.6 keV) X-ray surface brightness. While this approach offers advantages in terms of computational efficiency, it may be less capable of resolving subtle structures in dynamically complex regions. In contrast, the EP-FXT analysis employs adaptive binning with a SNR of 60 for spectral fitting, yielding higher spatial resolution maps of temperature and normalization, which better reveal localized pressure features.
In the pseudo-entropy map, the northern region displays an overall low entropy distribution. A similar low-entropy feature was also identified by Hudaverdi et al. (2005). This indicates that the fan-shaped structure is the disturbed core of the cluster.
At the center, the U-shaped surface brightness deficit exhibits high temperatures and entropy, indicating that it is filled with hot, low-density gas—features reminiscent of X-ray cavities typically driven by AGN feedback. However, A3571 lacks the characteristic radio emission commonly observed in such cavities (Venturi et al. 2002). This absence suggests that AGN feedback is unlikely to be the dominant cause, and that alternative mechanisms—such as mergers or gas sloshing—offer a more plausible explanation for the origin of the U-shaped structure.
4 DISCUSSION
Although the X-ray morphology of A3571 looks like relaxed, its residual map and 2D thermodynamic maps reveal disturbed structure in the inner region. Considering its galaxy density distribution is also elongated along the north–south direction. To identify possible merging signals, we select two sectors covering the brightness excesses in the northwest and southwest for detailed analysis.
Fig. 6 shows the sectors defined in the northwest and southwest directions, from which the radial surface brightness profiles were extracted and fitted using the best-fit model. In addition, spectra were extracted from each sector and fitted with a TBABS * APEC model, resulting in radial temperature profiles (see Fig. 8).
The results show that, in the northwestern direction, the radial surface brightness shows a deficiency within a radius of compared to the model, which corresponds to the outer boundary of the U-shaped structure in that direction. Between and , the surface brightness shows an excess corresponding to the fan-shaped structure in the residual map. Around , the radial temperature profile shows no pronounced discontinuity and exhibits only a slight increase. However, the line-of-sight projection of the cold sloshing tail could reduce the apparent temperature contrast. Alternatively, the presence of a weak cold front at this radius cannot be excluded.
In the southwestern direction, the surface brightness residual exhibits a deficiency within a radius of , while at approximately , the brightness shifts from an excess to a deficiency. This location may correspond to the outer boundary of the enhanced-brightness region in the southwest. However, within this region, the overall temperature remains consistently above 6 keV, but the radial temperature profile shows no pronounced discontinuity; consequently, there is no evidence of shock heating.
By combining the 1D (see Fig. 5) and 2D (see Fig. 7) temperature analyses, it is clear that the northern region is significantly cooler than the south. In Fig. 5, the temperature drop around is primarily associated with the northern sector, which corresponds in Fig. 7 to a fan-shaped region of low temperature and low entropy. In contrast, the XMM-Newton results do not show a pronounced temperature decrease, likely due to the limited FoV and reduced sensitivity. Nevertheless, the XMM-Newton data also reveal a temperature asymmetry between the north and south. We therefore suggest that this feature likely results from dynamical activity along the north–south axis.
Based on the above analysis, we propose that the structure of A3571 originates from gas sloshing triggered by the off-axis passage of a low-mass subcluster moving from south to north. The sloshing displaces low-entropy gas from the cool core, producing a fan-shaped brightness excess to the north. At the same time, the central region, originally filled with dense, cool gas, is replenished with diffuse hot gas, resulting in the observed U-shaped surface brightness deficit. This is consistent with the results of numerical simulations (Ascasibar and Markevitch 2006). These simulations show that even a low-mass subcluster, during an off-axis passage, can gravitationally disturb the main cluster, shifting its central mass peak and, through a reversal of ram pressure, eject cool gas from the bottom of the potential well. This process initiates persistent sloshing that can naturally generate fan-shaped distributions of cool gas and U-shaped surface brightness deficits similar to those observed in A3571.
5 Conclusion
We performed a systematic X-ray analysis of the galaxy cluster A3571, combining optical observations with X-ray data obtained from the EP-FXT telescope, which features a large FoV and low particle background. We measured the cluster’s outer properties beyond in imaging domain. The main results of this work are summarized as follows:
1.
Our data provides a high-quality image of A3571 within radius . It exhibits a relatively regular morphology. Both the surface brightness and temperature profiles exhibiting typical cool-core characteristics.
2.
Two prominent regions of surface brightness excesses are detected within of the cluster center, located to the north and southwest. The southwestern excess is associated with high-temperature features, whereas the northern excess corresponds to a cooler structure. The northern low-temperature feature is most likely produced by gas sloshing, occurring primarily along the line of sight and triggered by the off-axis passage of a low-mass subcluster.
3.
The temperature distribution of A3571 shows a pronounced north–south asymmetry, consistent with previous XMM-Newton observations. A similar north–south elongation is also evident in the optical galaxy density distribution, suggesting that the cluster’s merger activity likely occurred along this axis.
4.
We detect no thermal emission from the direction of A3572, nor do we find any excess in the density of optical galaxies; therefore, we conclude that this cluster does not exist. In contrast, the galaxy density distribution and X-ray morphology indicate that A3571 is still recovering from a minor merger and is currently in the post-merger phase.
The results of this study are limited by the calibration uncertainties of EP-FXT, which primarily affect measurements of the outer temperature and thermodynamic properties. Data from the formal observational phase (beginning in June 2024) are not affected by this issue. We expect that more comprehensive studies of the hot gas in galaxy clusters will be possible with these improved data.
Acknowledgements.
This work has been supported by the National Natural Science Foundation of China No. 12573003, the National Key Research and Development Program of China (No. 2023YFC2206704), National Key R&D Program of China No. 2025YFF0511104, the International Partnership Program of Chinese Academy of Sciences, Grant No. 013GJHZ2024015FN and the China Manned Space Program with grant No. CMS-CSST-2025-A04. A.L. acknowledges the supports from the National Natural Science Foundation of China (Grant No. 12588202). This work is based on data obtained with Einstein Probe, a space mission supported by Strategic Priority Program on Space Science of Chinese Academy of Sciences, in collaboration with ESA, MPE and CNES (Grant No. XDA15310000). EP is a space mission supported by Strategic Priority Program on Space Science of Chinese Academy of Sciences, in collaboration with ESA, MPE and CNES (Grant No.XDA15310303, No.XDA15310103, No.XDA15052100).References
- The Origin of Cold Fronts in the Cores of Relaxed Galaxy Clusters. ApJ 650 (1), pp. 102–127. External Links: Document, astro-ph/0603246, ADS entry Cited by: §1, §4.
- The Chemical Composition of the Sun. ARA&A 47 (1), pp. 481–522. External Links: Document, 0909.0948, ADS entry Cited by: §3.1.3.
- Overview of the DESI Legacy Imaging Surveys. AJ 157 (5), pp. 168. External Links: Document, 1804.08657, ADS entry Cited by: §2.2.
- An X-ray flux-limited sample of clusters of galaxies : evidence for evolution of the luminosity function.. MNRAS 245, pp. 559. External Links: ADS entry Cited by: §1.
- A very deep Chandra observation of the Perseus cluster: shocks, ripples and conduction. MNRAS 366 (2), pp. 417–428. External Links: Document, astro-ph/0510476, ADS entry Cited by: §1.
- A very deep Chandra observation of the Perseus cluster: shocks, ripples and conduction. MNRAS 366 (2), pp. 417–428. External Links: Document, astro-ph/0510476, ADS entry Cited by: §3.2.2.
- Metal distribution in sloshing galaxy clusters: the case of A496. A&A 570, pp. A117. External Links: Document, 1407.6814, ADS entry Cited by: §1.
- Study of the structure of abell 3571: An XMM-Newton view. Advances in Space Research 36 (4), pp. 643–649. External Links: Document, ADS entry Cited by: §1, §1, Figure 5, Figure 5, §3.1.3, §3.1.3, §3.2.1, §3.2.2, §3.2.2, §3.2.2.
- Sloshing and spiral structures breeding a putative radio mini-halo in the environment of a cool-core cluster, Abell 795. MNRAS 531 (4), pp. 4060–4069. External Links: Document, 2405.19750, ADS entry Cited by: §1.
- Sloshing cold fronts in galaxy cluster Abell 2566. New A 111, pp. 102253. External Links: Document, 2405.10475, ADS entry Cited by: §1.
- The 0.6 mpc halo of the cd galaxy (mcg 05–33–002) in the cluster abell 3571 revealed by the co-addition of eight schmidt plates. Monthly Notices of the Royal Astronomical Society 251 (1), pp. 10P–13P. Cited by: §1, §1.
- The Cluster-Merger Shock in 1E 0657-56. arXiv e-prints, pp. arXiv:0803.1803. External Links: Document, 0803.1803, ADS entry Cited by: §1.
- Formation of Galaxy Clusters. ARA&A 50, pp. 353–409. External Links: Document, 1205.5556, ADS entry Cited by: §1.
- Physical Properties of the X-Ray Gas as a Dynamical Diagnosis for Galaxy Clusters. In Tracing Cosmic Evolution with Clusters of Galaxies, pp. 117. External Links: ADS entry Cited by: §1.
- Brightest Cluster Galaxies at the Present Epoch. ApJ 797 (2), pp. 82. External Links: Document, 1407.2260, ADS entry Cited by: §2.2.
- Inside a Beehive: The Multiple Merging Processes in the Galaxy Cluster Abell 2142. ApJ 863 (1), pp. 102. External Links: Document, 1806.10864, ADS entry Cited by: §1.
- A Textbook Example of a Bow Shock in the Merging Galaxy Cluster 1E 0657-56. ApJ 567 (1), pp. L27–L31. External Links: Document, astro-ph/0110468, ADS entry Cited by: §1.
- Chandra Observation of Abell 2142: Survival of Dense Subcluster Cores in a Merger. ApJ 541 (2), pp. 542–549. External Links: Document, astro-ph/0001269, ADS entry Cited by: §1.
- The Temperature Structure of 30 Nearby Clusters Observed with ASCA: Similarity of Temperature Profiles. ApJ 503 (1), pp. 77–96. External Links: Document, astro-ph/9711289, ADS entry Cited by: §1.
- Shocks and cold fronts in galaxy clusters. Phys. Rep 443 (1), pp. 1–53. External Links: Document, astro-ph/0701821, ADS entry Cited by: §1, §1.
- Temperature and total mass profiles of the A3571 cluster of galaxies. A&A 369, pp. 459–466. External Links: Document, astro-ph/0101412, ADS entry Cited by: §1.
- X-Ray Total Mass Estimate for the Nearby Relaxed Cluster A3571. ApJ 536 (1), pp. 73–78. External Links: Document, astro-ph/0001162, ADS entry Cited by: §1, §1.
- Substructure in the Cold Front Cluster Abell 3667. ApJ 693 (1), pp. 901–913. External Links: Document, 0811.3031, ADS entry Cited by: §1.
- Planck 2018 results. VI. Cosmological parameters. A&A 641, pp. A6. External Links: Document, 1807.06209, ADS entry Cited by: §1.
- Spectroscopic observations of the galaxy cluster a: 3571 sc: 1344-325. Astronomy and Astrophysics Supplement, Vol. 101, NO. 3/NOVI, P. 475, 1993 101, pp. 475. Cited by: §1.
- On the detection of multiple shock fronts in A1914 using deep Chandra X-ray observations. MNRAS 509 (4), pp. 5821–5835. External Links: Document, 2110.12297, ADS entry Cited by: §1.
- X-ray and optical observations of the Shapley supercluster in Hydra-Centaurus.. MNRAS 248, pp. 101. External Links: Document, ADS entry Cited by: §1.
- Gas sloshing, cold front formation and metal redistribution: the Virgo cluster as a quantitative test case. MNRAS 413 (3), pp. 2057–2077. External Links: Document, 1007.4209, ADS entry Cited by: §1.
- Cool core remnants in galaxy clusters. A&A 510, pp. A83. External Links: Document, 0910.4900, ADS entry Cited by: §1.
- X-ray observations of a subhalo associated with the NGC 4839 group infalling toward the Coma cluster. PASJ 68 (5), pp. 85. External Links: Document, 1607.07554, ADS entry Cited by: §3.2.2.
- Cool-core, X-ray cavities, and cold front revealed in RXCJ0352.9+1941 cluster by Chandra and GMRT observations. Journal of Astrophysics and Astronomy 45 (2), pp. 23. External Links: Document, 2404.19549, ADS entry Cited by: §1.
- A Compilation of Redshifts and Velocity Dispersions for ACO Clusters. ApJS 125 (1), pp. 35–71. External Links: Document, ADS entry Cited by: §2.2.
- Radio emission from the A3571 cluster complex: The final stage of a cluster merger?. A&A 385, pp. 39–54. External Links: Document, astro-ph/0201532, ADS entry Cited by: §1, §3.2.2.
- Clusters of galaxies. Physics Uspekhi 57 (4), pp. 317–341. External Links: Document, ADS entry Cited by: §1.
- A Moving Cold Front in the Intergalactic Medium of A3667. ApJ 551 (1), pp. 160–171. External Links: Document, astro-ph/0008496, ADS entry Cited by: §1.
- Is there a giant Kelvin-Helmholtz instability in the sloshing cold front of the Perseus cluster?. MNRAS 468 (2), pp. 2506–2516. External Links: Document, 1705.00011, ADS entry Cited by: §1.
- Substructure and dynamical state of 2092 rich clusters of galaxies derived from photometric data. MNRAS 436 (1), pp. 275–293. External Links: Document, 1307.0568, ADS entry Cited by: §2.2.
- Calibration of x-ray absorption in our galaxy. Monthly Notices of the Royal Astronomical Society 431 (1), pp. 394–404. Cited by: §3.1.3.
- Science objectives of the Einstein Probe mission. arXiv e-prints, pp. arXiv:2501.07362. External Links: Document, 2501.07362, ADS entry Cited by: §1.
- The Einstein Probe Mission. In Handbook of X-ray and Gamma-ray Astrophysics, C. Bambi and A. Sangangelo (Eds.), pp. 86. External Links: Document, ADS entry Cited by: §1.
- The Instrumental Background of EP/FXT. External Links: 2507.00510, Link Cited by: §2.1.
- Estimate of the background and sensitivity of the follow-up X-ray telescope onboard Einstein Probe. Astroparticle Physics 137, pp. 102668. External Links: Document, 2111.12976, ADS entry Cited by: §3.1.1.
- Data reduction and processing for the Follow-up X-ray Telescope onboard Einstein Probe. Radiation Detection Technology and Methods. External Links: Document, ADS entry Cited by: §2.1.
- Imaging-spectroscopy diagnosis of the giant sloshing spiral in the Virgo cluster with the Einstein Probe Follow-up X-ray Telescope. A&A 700, pp. A248. External Links: Document, 2507.07412, ADS entry Cited by: §1.
- Stirring Up the Pot: Can Cooling Flows in Galaxy Clusters be Quenched by Gas Sloshing?. ApJ 717 (2), pp. 908–928. External Links: Document, 0912.0237, ADS entry Cited by: §1.
Appendix A Comparison of temperatures
To evaluate the consistency between FXT-A and FXT-B, we performed spectral fits in the 1.0–8.0 keV band within the Voronoi-defined regions (see Sect.3.2.2), using the TBABS*APEC model with the same parameter settings as in Sect.3.1.3. The temperature measurements were then compared (see Fig. 9). The results show that the temperatures obtained with FXT-A and FXT-B are consistent within the uncertainties. Although FXT-A yields slightly higher values than FXT-B, this discrepancy is likely due to FXT-B’s larger effective area at low energies and calibration differences between the two instruments.
In addition, the observation of A3571 was carried out during the PV phase in March, whereas the formal calibration campaign was not systematically initiated until April. Consequently, in the absence of finalized calibration products, these data are unavoidably affected by calibration-related systematic uncertainties.